OpenStudy (anonymous):

Jose is paying for a $10 meal using bills in his wallet. He has five $1 bills and four $5 bills. If he selects two bills at random, one at a time from his wallet, without replacement, what is the probability that he will choose two $5 bills to pay for the meal? Show your work.

OpenStudy (anonymous):

\[\frac{4}{9}\times \frac{3}{8}\]

OpenStudy (anonymous):

then reduce to get \(\frac{1}{6}\)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:)